OpenStudy (anonymous):

Please Help! The equation of the line with slope 1/2 and passing through (3,7) Is a.) y=1/2x+7 b.) 7x+3y=2 c.)x-2y+11=0 d.) 2x+y=-7

OpenStudy (midhun.madhu1987):

The general form of line with Slope = m and passing through the points (x1, y1) is y - y1 = m(x - x1) Just plug in the values and check which option is correct...

OpenStudy (anonymous):

y - y1 = m(x - x1) y-7=1/2(x-3) y-7=1/2x-3/2 y=1/2x+11/2 I keep getting this answer but its not in the multiple choice answers am i doing something wrong?

OpenStudy (triciaal):

you are not wrong just need to change format to match choices y=1/2x+11/2 multiply everything by 2 get rid of fraction rearrange to put everything on one side set = 0 to you see option c now?

OpenStudy (anonymous):

Alright thanks! =)